# Physics Informed Generative Machine Learning for Accelerated Quantum-centric Supercomputing

### v2

@@ Abstract
Quantum centric supercomputing (QCSC) framework, such as sample-based quantum diagonalization (SQD) holds immense promise toward achieving practical quantum utility to solve challenging problems. QCSC leverages quantum computers to perform the classically intractable task of sampling the dominant fermionic configurations from the Hilbert space that have substantial support to a target state, followed by Hamiltonian diagonalization on a classical processor. However, noisy quantum hardware produces erroneous samples upon measurements, making robust and efficient configuration-recovery strategies essential for a scalable QCSC pipeline. Toward this, in this work, we introduce PIGen-SQD, an efficiently designed QCSC workflow that utilizes the capability of generative machine learning (ML) along with physics-informed configuration screening via implicit low-rank tensor decompositions for accurate fermionic state reconstruction. The physics-informed pruning is based on a class of efficient perturbative measures that, in conjunction with hardware samples, provide a substantial overlap with the target state. This distribution induces an anchoring effect on the generative ML models to stochastically explore only the dominant sector of the Hilbert space for effective identification of additional important configurations in a self-consistent manner. Our numerical experiments performed on IBM Heron R2 quantum processors demonstrate this synergistic workflow produces compact, high-fidelity subspaces that substantially reduce diagonalization cost while maintaining chemical accuracy under strong electronic correlations. By embedding classical many body intuitions directly into the generative ML model, PIGen-SQD advances the robustness and scalability of QCSC algorithms, offering a promising pathway toward chemically reliable quantum simulations on utility-scale quantum hardware.
@@ I Introduction
Quantum computing holds immense promise for efficiently solving complex and classically intractable problems. A key application area is quantum chemistry, which seeks to determine the ground and excited state energies of many-body fermionic systems by solving the time-independent Schrödinger equation within the Born-Oppenheimer approximation. Designing efficient and accurate quantum algorithms is key toward solving some of the open problems in chemistry such as computation of ligand binding affinity of a drug molecule at a given protein target site[1, 2, 3], accurate electronic structure calculations of the Fe-S clusters for a better understanding of biological nitrogen fixation under ambient conditions[4], or designing high temperature superconducting materials, which involve strong many-body correlation effects driven by Cu 3d electron bands[5]. However, with the current degree of hardware noise, quantum algorithms like the quantum phase estimation (QPE)[6, 7] is not viable due to its requirement of deep circuits and long coherence time. Consequently, near-term research focuses on hybrid quantum-classical approaches like the variational quantum eigensolver (VQE)[8, 9, 10], quantum approximate optimization algorithm (QAOA)[11] and projective quantum eigensolver (PQE)[12] where quantum processors are dedicated to the evaluation of expectation values and classical processors perform parameter optimization. However, optimization of such hybrid algorithms for practically relevant systems is not guaranteed due to Barren Plateaus[13, 14, 15] and far-from-optimal local traps[16] in the optimization landscape, prolonged optimization runtimes and steep measurement overhead. Reducing the resource requirements for achieving practical quantum advantage of such hybrid algorithms is still an active area of research[17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Despite such intensive research efforts, scaling up electronic structure calculations solely with such variational optimization algorithms remains a nontrivial challenge. In this scenario, alternative algorithmic developments are required to scale up quantum chemical calculations using quantum computers for practical quantum advantage.
@@
Toward this, recently a distinct class of hybrid architecture termed Quantum-Centric Supercomputing (QCSC) has garnered attention due to its enhanced noise resilience and capability of sampling classically hard probability distributions, presenting a potential path forward for larger-scale simulations. QCSC is an emerging paradigm characterized by the seamless integration of quantum computing resources with traditional High-Performance Computing (HPC) infrastructures. This unified architecture is designed to address highly complex, real-world computational problems by distributing the workload to optimally leverage the distinct advantages of both quantum and classical processors. Some of the recently developed QCSC algorithms are inspired by classical selected configuration interaction (SCI)[37, 38, 39, 40, 41] such as quantum SCI (QSCI)[42], sample-based quantum diagonalization (SQD)[43, 44]. In such QCSC methods, generally a state is prepared on a quantum computer that approximately represents the target state and repeated measurements are performed to obtain samples that produce a set of computational basis states. For a molecular system these sampled computational basis states represent important electronic configurations which have dominant contribution to the target electronic state. HPC is required post-sampling for projection of the Hamiltonian in this electronic configuration subspace and diagonalization to get the energy of the approximate state. However, due to the presence of hardware noise, the sampled distributions often contain a substantial number of states that violate fundamental physical symmetries of the molecular system, such as particle number or spin conservation, leading to erroneous results upon diagonalization. To filter these noise-induced configurations and faithfully recover the target state, SQD employs a configuration recovery (CR) module. CR expands the subspace to generate more dominant electronic configurations in a self-consistent manner based on the average occupancy of the corresponding spin-orbitals with subsequent diagonalization. Since sampling dominant configurations from Hilbert space is a computationally challenging task, SQD has the potential to provide advantage over classical SCI methods, provided it produces accurate energy estimations using comparatively reduced dimensional configuration subspace for diagonalization. While SQD with its configuration recovery can produce accurate results for chemically complex systems under strong electronic correlations[43], the accuracy and efficacy depend on the prepared state and the quality of the hardware samples which should have sufficient overlap with the target state for any practical advantage against classical methods. For this, chemistry-inspired hardware-efficient ansatz can be used with parameters optimized in a sub-optimal manner via VQE.[45, 46] However, such an approach adds a heuristic component to the workflow which does not guarantee a substantial overlap with the target state to show any deterministic practical advantage. To this end, recently active research is going on to enhance the performance of SQD to push its limits toward better sampling of the configurations[47, 48]. Given that the state fidelity and overall performance of the SQD pipeline is directly dependent on the quality of this recovery step, the development of a more robust and efficient configuration protocol with better convergence guarantees is critical for the optimal utilization of HPC resources and the scalability of QCSC algorithms as a whole.
@@ II Brief Overview of Quantum Centric Supercomputing
where are the composite hole-particle indices. The dimension of space for closed-shell systems is given by with () being the number of electrons in the up (down) spin sectors. Generally the FCI coefficients for some low-lying eigen states are obtained via projection and diagonalization of the Hamiltonian operator in the symmetry space using efficient diagonalization methods like Davidson’s algorithm[55]. However, since the dimension of the symmetry space grows combinatorially with the system size, FCI method cannot be applied to practically relevant systems. Usually space is sparse since a relatively huge number of configurations have zero coefficients and are redundant in the CI expansion Eq. (3). The subspace which contains determinants with only non-zero coefficients is to be referred to as the core space with number of determinants which is in practice much smaller than . Finding out the exact core space is an extremely non-trivial task and the holy grail of any CI method as the prior information of the core space drastically reduces the dimensionality of the diagonalization task for better scalability. Toward this, several selected CI methods have been developed recently for selecting subspaces that approximate core space while retaining the desired accuracy[37, 39, 40, 41, 56].
Quantum centric supercomputing (QCSC) or quantum selected CI (QSCI) methods are a class of hybrid quantum-classical algorithms designed to outsource this classically nontrivial core space selection task to a quantum processor. This is followed by projection and diagonalization of the corresponding Hamiltonian in the selected space on a classical computer for energy estimation. Toward this, the fermionic electronic-structure problem is mapped to the quantum-computing framework by encoding the M molecular spin-orbitals onto M qubits, such that each electronic configuration corresponds to a computational-basis state where . The single-qubit states or represents unoccupied and occupied spin-orbitals, respectively. Jordan-Wigner transformation is used to map the corresponding second-quantized fermionic operators to qubit operators. In the QCSC framework, initially a trial state on qubits (which represent molecular spin orbitals) is prepared on a quantum computer with a parameterized quantum circuit (PQC) that should have overlap with the true target state. This is ensured by preparing a state with a physically motivated parameterized ansatz structure of the general form where the unitary operator (ensured by the anti-Hermiticity of the operator ) explicitly introduces electronic corrleation by acting upon the reference state. This PQC is executed on a quantum computer to perform repeated measurements in the computational basis that produce a distribution of measurement outcomes
@@ III.1 Anchoring the Hardware Samples to Ground State via Many-Body Theories
Starting from the first principle many-body theory, an unbiased ensemble of the ground state configurations may be built up via Møller-Plesset perturbation theory (MPPT)[57, 58] with the perturbing potential where denotes the Hartree-Fock potential. The corresponding Hilbert space is split into a model space spanned by a reference function and a perturbative orthogonal space spanned by excited determinants defined by . In the perturbative expansion, the exact ground state can be written as
@@ III.2 Generative Machine Learning Based Fermionic State Reconstruction
where, is the partition function. Since explicit calculation of the partition function becomes computationally intractable, RBMs are usually trained on a set of vectors distributed according to some unknown probability distribution using the contrastive divergence (CD)[61] method which approximates the gradient of the log likelihood to learn the underlying distribution. RBM alternatively updates hidden and visible layers using conditional activation probabilities:
@@ IV Computational Details
All the numerical results in this work are obtained by in-house python based codes. MBPT calculations are done using two-body integrals and MP2 amplitudes obtained from PySCF[67] driver in Qiskit[68]. Three strongly correlated molecular systems , and in 6-31G basis with frozen core approximation involving 24, 32 and 40 spin orbitals correlating 8, 10 and 10 electrons respectively are studied in this work. The FCI, CCSD, CI with singles and doubles (CISD), MP2 and HF calculations were performed using PySCF with the frozen core approximation. Since FCI calculations are not feasible for , heat-bath configuration interaction (HCI)[39] calculations were done using PySCF to get reference energy values for accuracy comparison. The entire workflow involves multiple independent computational components orchestrated together as discussed below:
@@
Quantum state preparation and hardware sample generation: State preparation on a quantum hardware is done using the LUCJ ansatz [45] (Eq. (6)) with CCSD parameters used for jastrow decomposition. No further optimization of the LUCJ parameters is performed. For 24 qubits experiment 250000 shots were used while both 32 and 40 qubit experiments were performed using shots. The circuits were created using qiskit-addon-sqd, which uses the ffsim[69] package to generate the LUCJ circuits. For each circuit, we start with some generic linear layout and invoke the VF2PostLayout transpiler pass to find an isomorphic zig-zag layout (which is required for LUCJ circuits) that conforms to the lowest noise profile as obtained from the backend noise information. The LUCJ circuit is placed on this layout using the qiskit transpiler. We executed all the circuits on ibm_kingston, which is a 156 qubits IBM Heron R2 processor. The layouts used for the circuits corresponding to each of the molecules are shown in Fig. 1.
@@
Symbolic Sparse Tensor Operation toward Configuration Selection:
@@
To obtain the physics-informed support to ground state via perturbative measures, initially available molecular integrals (in molecular orbital basis) are filtered from the original sparse integral tensor using a threshold . All surviving integral values, together with their associated index combinations, are collected and arranged into an array to construct following the integral index pattern (). The computational cost of triples selection step is governed by the most expensive term which can be performed efficiently using the symbolic index operation technique discussed in Sec.III. For this efficient generation we combine all unique outer indices and having only one index common between them (see Appendix B) with the constraints and respectively, stemming from pruned and . Such an operation does not require explicit summations over contractible indices, and just one common index suffices the inclusion of the corresponding outer (non-contractible) indices. At this stage, it is important to note that the hierarchical process of configuration selection through perturbation theory automatically ensures that the lower order terms, double excitations () from the HF reference, are already pruned which implies . This provides the flexibility to avoid the prohibitive cost of explicit tensor contraction, saving huge number of operations to be performed. A comparison of the number of operations required for the pruned elements in the numerical worst case scenario for triples and quadruples selection are shown in Fig.2 against the formal asymptotic scaling of the term , CCSD and CCSDT. Fig. 2 shows the numerical limiting case of symbolic tensor manipulation with pruned elements for triples selection requires almost two orders-of-magnitude less number of operations compared to the asymptotic cost of explicit computations. The actual implementational cost for the same using the symbolic technique requires much less number of operations than this in all practical cases. A similar approach can be followed for constructing perturbative quadruples as well by breaking the associated contraction into intermediates and performing the index-based symbolic calculations. The algebraic terms calculated for selection of quadruply excited configurations are given in Appendix A. The corresponding numerical limiting case scales almost similar to CCSD as demonstrated in Fig.2.
@@
While this way of selection of important high-rank excited configurations through their lowest order perturbative measures is not theoretically exact from a conventional many-body viewpoint, we argue that an absolute accuracy (of choosing the important configurations) is not also warranted for; rather, this measure gives us a strong additional support to the hardware generated configuration samples. Dominant determinants that are missed out (due to noise or perturbatively) would eventually get picked up through generative ML. Also, in a similar spirit, any determinant which are spuriously picked up would get dropped (blacklisted) during the iterative fermionic state reconstruction as will be shown in numerical demonstrations.

### v3

# Physics-Informed Generative Machine Learning for Accelerated Quantum-Centric Supercomputing

@@ Abstract
Quantum centric supercomputing (QCSC) framework, such as sample-based quantum diagonalization (SQD) holds immense promise toward achieving practical quantum utility to solve classically hard sampling problems in quantum chemistry. QCSC leverages quantum computers to perform the classically intractable task of sampling the dominant fermionic configurations from the Hilbert space that have substantial support to a target state, followed by Hamiltonian diagonalization on a classical processor. However, noisy quantum hardware produces erroneous samples upon measurements, making robust and efficient configuration-recovery strategies essential for a scalable QCSC pipeline. Toward this, in this work, we introduce PIGen-SQD, an efficiently designed QCSC workflow that utilizes the capability of generative machine learning (ML) along with physics-informed configuration screening via implicit low-rank tensor decompositions for accurate fermionic state reconstruction. The physics-informed pruning is based on a class of efficient perturbative measures that, in conjunction with samples drawn from quantum hardware, provide a substantial overlap with the target state. This distribution induces an anchoring effect on the generative ML models to stochastically explore only the dominant sector of the Hilbert space for effective identification of additional important configurations in a self-consistent manner. Our numerical experiments performed on IBM Heron R2 and R3 quantum processors with up to 52 qubit experiments demonstrate this synergistic workflow produces compact, high-fidelity subspaces that substantially reduce diagonalization cost while maintaining chemical accuracy under strong electronic correlations. With such a concerted integration of classical many-body intuitions, generative ML and quantum computers as a sampling engine, PIGen-SQD offers a promising pathway toward accurate and systematically improvable quantum simulations on utility-scale quantum hardware.
@@ I Introduction
Quantum computing holds immense promise for efficiently solving complex and classically intractable problems. A key application area is quantum chemistry, which seeks to determine the ground and excited state energies of many-body fermionic systems by solving the time-independent Schrödinger equation within the Born-Oppenheimer approximation. Designing efficient and accurate quantum algorithms is key toward solving some of the open problems in chemistry such as computation of ligand binding affinity of a drug molecule at a given protein target site[1, 2, 3], accurate electronic structure calculations of the Fe-S clusters for a better understanding of biological nitrogen fixation under ambient conditions[4], or designing high temperature superconducting materials, which involve strong many-body correlation effects driven by Cu 3d electron bands[5]. However, with the current degree of hardware noise, quantum algorithms like the quantum phase estimation (QPE)[6, 7] is not viable due to its requirement of deep circuits and long coherence time. Consequently, near-term research focuses on hybrid quantum-classical approaches like the variational quantum eigensolver (VQE)[8, 9, 10], quantum approximate optimization algorithm (QAOA)[11] and projective quantum eigensolver (PQE)[12] where quantum processors are dedicated to the evaluation of expectation values and classical processors perform parameter optimization. However, optimization of such hybrid algorithms for practically relevant systems is not guaranteed due to Barren Plateaus[13, 14, 15] and far-from-optimal local traps[16] in the optimization landscape, prolonged optimization runtimes and steep measurement overhead. Reducing the resource requirements for achieving practical quantum advantage of such hybrid algorithms is still an active area of research[17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Despite such intensive research efforts, scaling up electronic structure calculations solely with such variational optimization algorithms remains a nontrivial challenge. In this scenario, alternative algorithmic developments are required to scale up quantum chemical calculations using quantum computers for practical quantum advantage.
@@
Toward this, recently a distinct class of hybrid architecture termed Quantum-Centric Supercomputing (QCSC) has garnered attention due to its enhanced noise resilience and capability of sampling classically hard probability distributions, presenting a potential path forward for larger-scale simulations. QCSC is an emerging paradigm characterized by the seamless integration of quantum computing resources with traditional High-Performance Computing (HPC) infrastructures. This unified architecture is designed to address highly complex, real-world computational problems by distributing the workload to optimally leverage the distinct advantages of both quantum and classical processors. Some of the recently developed QCSC algorithms are inspired by classical selected configuration interaction (SCI)[40, 41, 42, 43, 44] such as quantum SCI (QSCI)[45], sample-based quantum diagonalization (SQD)[46, 47] and its different variants[45, 46, 48, 47, 49, 50, 51, 52]. In such QCSC methods, generally a state is prepared on a quantum computer that approximately represents the target state and repeated measurements are performed to obtain samples that produce a set of computational basis states. For a molecular system these sampled computational basis states represent important electronic configurations which have dominant contribution to the target electronic state. HPC is required post-sampling for projection of the Hamiltonian in this electronic configuration subspace and diagonalization to get the energy of the approximate state. However, due to the presence of hardware noise, the sampled distributions often contain a substantial number of states that violate fundamental physical symmetries of the molecular system, such as particle number or spin conservation, leading to erroneous results upon diagonalization. To filter these noise-induced configurations and faithfully recover the target state, SQD employs a configuration recovery (CR) module. CR expands the subspace to generate more dominant electronic configurations in a self-consistent manner based on the average occupancy of the corresponding spin-orbitals with subsequent diagonalization. Since sampling dominant configurations from Hilbert space is a computationally challenging task, SQD has the potential to provide advantage over classical SCI methods, provided it produces accurate energy estimations using comparatively reduced dimensional configuration subspace for diagonalization. While SQD with its configuration recovery can produce accurate results for chemically complex systems under strong electronic correlations[46], the accuracy and efficacy depend on the prepared state and the quality of the hardware samples which should have sufficient overlap with the target state for any practical advantage against classical methods. For this, chemistry-inspired hardware-efficient ansatz can be used with parameters optimized in a sub-optimal manner via VQE.[53, 54] However, such an approach adds a heuristic component to the workflow which does not guarantee a substantial overlap with the target state to show any deterministic practical advantage. To this end, recently active research is going on to enhance the performance of SQD to push its limits toward better sampling of the configurations[50, 51]. Given that the state fidelity and overall performance of the SQD pipeline is directly dependent on the quality of this recovery step, the development of a more robust and efficient configuration recovery protocol with better convergence guarantees is critical for the optimal utilization of HPC resources and the scalability of QCSC algorithms as a whole.
@@ II Brief Overview of Quantum Centric Supercomputing
where are the composite hole-particle indices. The dimension of space for closed-shell systems is given by
with () being the number of electrons in the up (down) spin sectors. Generally the FCI coefficients for some low-lying eigen states are obtained via projection and diagonalization of the Hamiltonian operator in the symmetry space using efficient diagonalization methods like Davidson’s algorithm[61]. However, since the dimension of the symmetry space grows combinatorially with the system size, FCI method cannot be applied to practically relevant systems. Usually space is sparse since a relatively huge number of configurations have zero coefficients and are redundant in the CI expansion Eq. (3). The subspace which contains determinants with only non-zero coefficients is to be referred to as the core space with number of determinants which is in practice much smaller than . Finding out the exact core space is an extremely non-trivial task and the holy grail of any CI method as the prior information of the core space drastically reduces the dimensionality of the diagonalization task for better scalability. Toward this, several selected CI methods have been developed recently for selecting subspaces that approximate core space while retaining the desired accuracy[40, 42, 43, 44, 62].
Quantum centric supercomputing (QCSC) or quantum selected CI (QSCI) methods are a class of hybrid quantum-classical algorithms designed to outsource this classically nontrivial core space selection task to a quantum processor. This is followed by projection and diagonalization of the corresponding Hamiltonian in the selected space on a classical computer for energy estimation. Toward this, the fermionic electronic-structure problem is mapped to the quantum-computing framework by encoding the M molecular spin-orbitals onto M qubits, such that each electronic configuration corresponds to a computational-basis state where . The single-qubit states or represents unoccupied and occupied spin-orbitals, respectively. Jordan-Wigner transformation is used to map the corresponding second-quantized fermionic operators to qubit operators. In the QCSC framework, initially a trial state on qubits (which represent molecular spin orbitals) is prepared on a quantum computer with a parameterized quantum circuit (PQC) that should have overlap with the true target state. This is ensured by preparing a state with a physically motivated parameterized ansatz structure of the general form
where the unitary operator (ensured by the anti-Hermiticity of the operator ) explicitly introduces electronic corrleation by acting upon the reference state. This PQC is executed on a quantum computer to perform repeated measurements in the computational basis that produce a distribution of measurement outcomes
@@ III.1 Anchoring the Hardware Samples to Ground State via Many-Body Theories
Starting from the first principle many-body theory, an unbiased ensemble of the ground state configurations may be built up via Møller-Plesset perturbation theory (MPPT)[63, 64] with the perturbing potential
where denotes the Hartree-Fock potential. The corresponding Hilbert space is split into a model space spanned by a reference function and a perturbative orthogonal space spanned by excited determinants defined by
In the perturbative expansion, the exact ground state can be written as
@@ III.2 Generative Machine Learning Based Fermionic State Reconstruction
where,
is the partition function. Since explicit calculation of the partition function becomes computationally intractable, RBMs are usually trained on a set of vectors distributed according to some unknown probability distribution using the contrastive divergence (CD)[67] method which approximates the gradient of the log likelihood to learn the underlying distribution. RBM alternatively updates hidden and visible layers using conditional activation probabilities:
@@ IV Computational Details
All the numerical results in this work are obtained by in-house python based codes. MBPT calculations are done using two-body integrals and MP2 amplitudes obtained from PySCF[73] driver in Qiskit[74]. Three strongly correlated molecular systems, (equilibrium O-H bond distance , bond angle ), () and () are studied for different geometries using 6-31G basis set with frozen core approximation involving 24, 32 and 40 spinorbitals correlating 8, 10 and 10 electrons respectively. The hardware mapping of the corresponding LUCJ ansatz on IBM Heron R2 processor requires 27, 36 and 45 qubits respectively that includes the qubits associated with Jordan-Wigner mapped spin orbitals and ancilla qubits. Numerical experiments are also performed using IBM Heron R3 processor on in the cc-pVDZ basis, which comprises 46 spin orbitals, that requires a total of 52 qubits. Details of the quantum hardware layout and the qubit assignments for all the numerical experiments are given in Appendix D. The FCI, CCSD, CI with singles and doubles (CISD), MP2 and HF calculations were performed using PySCF with the frozen core approximation. Since FCI calculations are not feasible for (6-31G basis) and (cc-pVDZ basis) under consideration, heat-bath configuration interaction (HCI)[42] calculations were done using DICE-solver[44] to get reference energy values for accuracy comparison. The entire workflow involves multiple independent computational components orchestrated together as discussed below:
@@
Quantum state preparation and hardware sample generation: State preparation on a quantum hardware is done using the LUCJ ansatz (Eq. (8)) with CCSD parameters used for jastrow decomposition. No further optimization of the LUCJ parameters is performed. For 27 qubits experiment 250000 shots were used while 36, 45 and 52 qubit experiments were performed using a fixed shots due to limited availability of quantum processor time and shot budget. Ideally, at least (see (4)) number of shots are required for an efficient and meaningful sampling task from a quantum hardware. However, in our case only for the study of in 6-31G basis () the number of shots is comparable to the dimension of the symmetry space. The circuits were created using qiskit-addon-sqd, which uses the ffsim[75] package to generate the LUCJ circuits. For each circuit, we start with some generic linear layout and invoke the VF2PostLayout transpiler pass to find an isomorphic zig-zag layout (which is required for LUCJ circuits) that conforms to the lowest noise profile as obtained from the backend noise information. The LUCJ circuit is placed on this layout using the qiskit transpiler. We executed the circuits on ibm_kingston and ibm_boston, having 156 qubits IBM Heron R2 and R3 processors respectively. The layouts used for the circuits corresponding to each of the molecules are shown in Appendix D (Fig. 10 and 11).
@@
Numerical Worst-Case Scenario for Symbolic Sparse Tensor Operation toward Configuration Selection:
@@
To obtain the physics-informed support to ground state via perturbative measures, initially available molecular integrals (in molecular orbital basis) are filtered from the original sparse integral tensor using a threshold . All surviving integral values, together with their associated index combinations, are collected and arranged into an array to construct following the integral index pattern (). The most expensive computational step for triples selection is governed by the term which can be performed efficiently using the symbolic index operation technique discussed in Sec.III. For this efficient generation we combine all unique outer indices and having only one index common between them (see Appendix B) with the constraints and respectively, stemming from pruned and . Such an operation does not require explicit summations over contractible indices, and just one common index suffices the inclusion of the corresponding outer (non-contractible) indices. At this stage, it is important to note that the hierarchical process of configuration selection through perturbation theory automatically ensures that the lower order terms, double excitations () from the HF reference, are already pruned which implies . This provides the flexibility to avoid the prohibitive cost of explicit tensor contraction, saving huge number of operations to be performed. A comparison of the number of operations required for the pruned elements in the numerical worst case scenario for triples and quadruples selection are shown in Fig.1 against the formal asymptotic scaling of the term , CCSD and CCSDT. Fig. 1 shows the numerical limiting case of symbolic tensor manipulation with pruned elements for triples selection requires almost two orders-of-magnitude less number of operations compared to the asymptotic cost of explicit computations. The actual implementational cost for the same using the symbolic technique requires much less number of operations than this in all practical cases. A similar approach can be followed for constructing perturbative quadruples as well by breaking the associated contraction into intermediates and performing the index-based symbolic calculations. The algebraic terms calculated for selection of quadruply excited configurations are given in Appendix A. The corresponding numerical limiting case for quadruples selection scales almost similar (or marginally higher) to CCSD as demonstrated in Fig.1. Here, we have considered a selection threshold which can be tuned to a larger value to reduce the number of operations further for sparse tensor contractions that would miss some of the configurations (with comparatively smaller coefficients in the expansion).
@@
While this way of selection of important high-rank excited configurations through their lowest order perturbative measures is not theoretically exact from a conventional many-body viewpoint, we argue that an absolute accuracy (of choosing the important configurations) is not also warranted for; rather, this measure gives us a strong additional support to the hardware generated configuration samples. Dominant determinants that are missed out (due to noise or perturbatively) would eventually get picked up through generative ML. Also, in a similar spirit, any determinant which are spuriously picked up would get dropped (blacklisted) during the iterative fermionic state reconstruction as will be shown in numerical demonstrations. Owing to this flexibility in the workflow, the symbolic tensor contractions can be restricted to terms with favorable scaling complexity, while still generating a substantial number of dominant configurations of a given rank that exhibit significant overlap with the target state. A detailed discussion regarding this is presented in Appendix B